OpenStudy (anonymous):

Write the equation of a parabola whose axis of sym. is x= -1. Write the equation in vertex form.

OpenStudy (anonymous):

vertex form is \[y=a(x-h)^2+k\] plug in \[x^2+2x+3\] find vertex which is \[\frac{ -b}{ 2(a) }\] which is x then plug it into \[(\frac{ -b }{ 2(a) })^2+2(\frac{ -b }{ 2(a) })+3\] (x,y) x=h y=k plug in a for \[ax^2+bx+c\] example of answer:\[y=2(x-5)+4\]